A new manufacturing method is supposed to increase the average life span of electronic com-
ponents, while the variance of the life span is expected to stay the same. Using the previous
manufacturing method, the average life span was 112.5 hours with a variance 12 hours. The man-
ufacturer wishes to establish the new average life span by measuring the life spans of a sample of
components manufactured using the new method.
1
Expert's answer
2021-09-07T18:56:55-0400

(a) (a) What sample size should be used, if the manufacturer wishes to establish the new average

life span to within 1 hours, with 90% level of confidence?

The critical value for "\alpha=0.1" is "z_c=z_{1-\alpha/2}=1.6449."

"z_c\times\dfrac{\sigma}{\sqrt{n}}\leq1""n\geq(\dfrac{z_c\sigma}{1})^2""n\geq(\dfrac{1.6449(12)}{1})^2""n\geq390"



(b)How will the required sample size change, if the manufacturer wishes to establish the new

average life span to within 1 hours, with 95% level of confidence?

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

"z_c\times\dfrac{\sigma}{\sqrt{n}}\leq1""n\geq(\dfrac{z_c\sigma}{1})^2""n\geq(\dfrac{1.96(12)}{1})^2""n\geq554"



(c) How will the required sample size change, if the manufacturer wishes to establish the new

average life span to within 1/2 hours, with 90% level of confidence?

The critical value for "\alpha=0.1" is "z_c=z_{1-\alpha/2}=1.6449."

"z_c\times\dfrac{\sigma}{\sqrt{n}}\leq1/2""n\geq(\dfrac{z_c\sigma}{1/2})^2""n\geq(\dfrac{1.6449(12)}{1/2})^2""n\geq1559"

The required sample size should be increased by 4 times to establish the new average life span to within 1/2 hours, with 90% level of confidence.